# Collaborative Multi-Modal Coding for High-Quality 3D Generation

### v1

@@ Abstract
Index Terms:
Image-to-3D Generation, Diffusion Model, Multi-modality
@@ I Introduction
Prior arts on 3D generation could be further categorized into three: 1) optimization-based methods [6] using Score Distillation Sampling (SDS); 2) reconstruction-based feedforward models [7, 8, 9] which employ a transformer to regress 3D representation given sparse-view images; 3) 3D-native feedforward models [10, 11, 12, 13, 14] with 3D diffusion or autoregressive framework. Current 3D generative approaches, despite architectural variations, predominantly adhere to monolithic modality paradigms—typically relying on colored renderings as main training sources. While RGB images provide dense texture priors of 3D assets (e.g., material reflectance, specular highlights), this photometric modality suffers from fundamental limitations: 1) geometric ambiguity in occluded regions, and 2) topological uncertainty due to projective viewpoint. These modality-specific constraints limit the scalability and quality of existing 3D generative models.
@@
To address this challenge, we propose TriMM, a collaborative multi-modal coding for high-quality feed-forward 3D generation. At the core of our framework is a unified latent coding that synergistically integrates photometric and geometric representations from multimodal data (RGB, RGBD, and point clouds), jointly benefiting from the high-frequency texture details encoded by RGB images and metric-accurate topologies encoded by point clouds and depth maps. Our framework employs dedicated modality-specific encoders coupled with a shared decoder architecture to map heterogeneous inputs into a unified triplane-structured latent representation. To effectively harness the complementary strengths of multimodal inputs, we introduce the reconstruction-based mechanism in the generative model that explicitly guides the model to discern and optimally utilize the distinctive advantages of each modality. To avoid distortion at the large elevation angle and raise the robustness and performance of multi-modal coding, we adopt auxiliary 2D (e.g., RGB, mask, depth loss) and 3D supervision (SDF loss). This collaborative multi-modal coding further facilitates efficient generative modeling which trains a lightweight latent diffusion model conditioned on input images. The well-structured latent representation space established by our multimodal framework enables streamlined generative model training, requiring neither intricate hyperparameter adjustments nor extensive optimization cycles. Extensive evaluations on standard benchmarks [4, 15] demonstrate that, by leveraging multi-modality information, TriMM achieves competitive performance with recent state-of-the-art methods in generating high-quality 3D assets from images, even when trained on small-scale datasets illustrate in Fig. 1. Beyond exploring the potential of current 3D data, our multimodal encoding architecture inherently supports the tokenization of real-world multi-modal inputs through its extensible design. This framework provides a novel and promising pathway to systematically address the challenge of 3D training data scarcity.
@@
  • We propose TriMM, a Collaborative Multi-Modal Coding that marries geometric and photometric information in multi-modal data into a unified space for high-quality 3D generation.
@@
  • Based on multi-modal coding, we build a generative framework that can leverage the potential and strength of multi-modal data while avoiding the weakness of specific modalities via a special reconstruction loss for 3D generation.
@@ II-B Feed-forward methods
II-B2 3D-native diffusion models:
As one of the powerful generative models, diffusion model has demonstrated its impressive potential in the 3D generation task. Point-E [24] proposes a diffusion model based on point cloud which is flexible but lacks the capability of representing watertight and solid surface of 3D content. To handle this, MeshDiffusion [25] and DiffRF [12] adopt mesh and voxel as the representation. Although explicit representation has superiority in querying and evaluating, those explicit 3D representation methods are memory-expensive when using high-resolution grids. Different from explicit representation, the neural field-based methods [26, 27] adopt an implicit way to encode the geometry and texture information. Despite low memory loading in high-resolution grids, it will cost more time to evaluate. Compared with pure explicit or implicit methods, the hybrid representation triplane achieves a promising trade-off between efficiency and performance [28, 29, 13, 14]. Recently, TRELLIS [30] proposes a powerful 3D generation framework that uses large-scale high-quality 3D data and 2D&3D reference input for reconstruction. Despite introducing multi-modal data, the early fuse framework limits its scalability across different modalities and still suffers from the scarcity of 3D data. To avoid this problem, we integrate the strength of multi-modal data in the later module. In this way, our model not only utilizes the potential of current 3D data to raise the performance but also inherently supports the tokenization of heterogeneous multi-modal inputs through its extensible design.
II-B3 Reconstruction-based methods:
@@
II-C Collaborative multi-modal coding
As the key component of our TriMM, Collaborative Multi-Modal Coding consists of 1) modality-specific encoders and 2) collaborative multi-modal decoders. In specific, the encoder part includes three branches, i.e., RGB, RGBD, and point cloud branch. To leverage dense texture information of RGB and spatial details of depth and point cloud, we apply three different encoders to project multi-modal data to the same latent space. Besides, since our Collaborative Multi-Modal Coding is trained on the public 3D dataset, it can be utilized to extend the scarce 3D data for generation tasks. By converting different RGB images, RGBD images, and point cloud datasets into a shared latent space spanned by triplane, we can introduce diverse training data to raise the capability of modeling complex structures.
@@ II-C Collaborative multi-modal coding
II-D Triplane latent diffusion model
@@
II-E Implementation details
@@
Training stages In this subsection, we report more details about our TriMM. We compare the parameters, training, and inference time of different modules in Table I. We use 24 NVIDIA A100 GPUs to train our model. Note that the inference time is tested on NVIDIA A100 GPUs. We split our model into two stages: 1) multi-modal coding training including RGB, RGBD, point cloud tokenizer, and shared decoder; 2) Triplane latent diffusion training including VAE and image-conditional diffusion model. For the 1st stage, we first train a RGB reconstruction model similar to other LRM-based methods, e.g., InstantMesh [9]. Besides, to accelerate the convergence of our multi-modal coding, we use the pre-trained RGB triplane to initialize other triplanes. After obtaining all multi-modality triplanes, we train a typical latent diffusion model (2nd stage). Considering the application, we adopt a single image as the condition to train the generative model.
@@
II-F Representation learning results
@@
II-G Triplane diffusion results
II-G1 Qualitative evaluations
In this subsection, we compared our method with our baseline and the most recently existing works including TripoSR [41], CRM [44], LGM [8], InstantMesh [9], and TRELLIS [30]. Baseline represents the generative model trained only on RGB triplanes. We adopt the pre-trained model LGM (trained on 80k data), CRM (trained on 376k data), InstantMesh (trained on 270k data), and TRELLIS (trained on 500k data). We compared our TriMM with other SOTA methods qualitatively in Fig. 6. We can notice that the generated 3D meshes of our TriMM are more stable when facing objects with different appearances and structures. TripoSR can generate a promising texture for 3D assets, especially in the frontal view. However, because triplane representation lacks the robust capability to express complex geometry structures, it lent to generate flattened 3D objects. Besides, since LGM, CRM, and InstantMesh are built based on a multi-view diffusion model, the performance of those methods is constrained by the performance of the diffusion model which decreases the final performance of image-to-3D. Relying on high-quality large-scale training data, TRELLIS achieves superior robustness and performance in various 3D assets. Different from integrating 2D and 3D data directly, we introduce the multi-modal coding module based on multi-modal data, thereby raising the robustness when facing complex structures and maintaining the capability of extension. Despite using smaller data, by exploring the potential of current 3D assets, our model achieves a competitive performance compared with TRELLIS. More qualitative results are shown in Fig. 8.
@@ II-H Ablation studies
II-H1 Analysis about multi-modal data.
@@
II-H2 Analysis about reconstruction loss.
@@ II-H3 Analysis of 2D/3D supervision
II-I Further analysis
II-I1 Multi-modal data to 3D.
@@ II-I Further analysis
II-I2 Introducing multi-modalities for 3D generation
@@
II-I4 Limitation and future work
Despite promising results, our method still has some limitations. The inefficient information utilization inherent in triplane representation imposes memory-bound restrictions on resolution enhancement, consequently diminishing their capacity for detailed feature characterization. Developing a new representation to improve the quality while maintaining a reasonable distribution is our next work. Besides, considering the different configurations of existing multi-modality data, it is still many bottlenecks in introducing multi-modality data for 3D modeling. For example, most of the point cloud dataset has only geometric information and most of the RGBD dataset focuses on scenes. The implementation of object segmentation masks, while effective for the localization of target objects, cannot inherently compensate for the resolution degradation in RGBD sensing data. We will carry them in our future work.

### v2

@@ I Introduction
Prior arts on 3D generation could be further categorized into three: 1) optimization-based methods [29] using Score Distillation Sampling (SDS); 2) reconstruction-based feedforward models [15, 36, 49] which employ a transformer to regress 3D representation given sparse-view images; 3) 3D-native feedforward models [14, 8, 26, 4, 5, 18, 7, 17, 25] with 3D diffusion or autoregressive framework. Current 3D generative approaches, despite architectural variations, predominantly adhere to monolithic modality paradigms—typically relying on colored renderings as main training sources. While RGB images provide dense texture priors of 3D assets (e.g., material reflectance, specular highlights), this photometric modality suffers from fundamental limitations: 1) geometric ambiguity in occluded regions, and 2) topological uncertainty due to projective viewpoint. These modality-specific constraints limit the scalability and quality of existing 3D generative models.
@@
To address this challenge, we propose TriMM, a collaborative multi-modal coding for high-quality feed-forward 3D generation. At the core of our framework is a unified latent coding that synergistically integrates photometric and geometric representations from multimodal data (RGB, RGBD, and point clouds), jointly benefiting from the high-frequency texture details encoded by RGB images and metric-accurate topologies encoded by point clouds and depth maps. Our framework employs dedicated modality-specific encoders coupled with a shared decoder architecture to map multi-modality information into a unified triplane-structured latent representation. To effectively harness the complementary strengths of multi-modal inputs, we introduce the reconstruction-based mechanism in the diffusion model that explicitly guides the model to discern and optimally utilize the distinctive advantages of each modality. To avoid distortion at the large elevation angle and raise the robustness and performance of multi-modal coding, we adopt auxiliary 2D (e.g., RGB, mask, depth loss) and 3D supervision (SDF loss). This collaborative multi-modal coding further facilitates efficient generative modeling which trains a lightweight latent diffusion model conditioned on input images. The well-structured latent representation space established by our multimodal framework enables streamlined generative model training, requiring neither intricate hyperparameter adjustments nor extensive optimization cycles. Extensive evaluations on standard benchmarks [10, 11] demonstrate that, by leveraging multi-modality information, TriMM achieves competitive performance with recent state-of-the-art methods in generating high-quality 3D assets from images, even when trained on small-scale datasets illustrate in Fig. 1. Beyond exploring the potential of current 3D data, our multimodal encoding architecture inherently supports the tokenization of real-world multi-modal inputs through its extensible design. This framework provides a novel and promising pathway to systematically address the challenge of 3D training data scarcity.
@@
  • We propose TriMM, a Collaborative Multi-Modal Coding framework that fuses geometric and photometric information from multiple modalities into a unified latent space for high-quality 3D generation, using modality-specific encoders and a shared decoder.
@@
  • Building on this multi-modal coding, we develop a generative framework that exploits the complementary strengths of different modalities while mitigating modality-specific weaknesses and the ambiguity through a tailored reconstruction loss for 3D generation.
@@ II-B Feed-forward methods
II-B2 3D-native diffusion models
As one of the powerful generative models, diffusion model has demonstrated its impressive potential in the 3D generation task. Point-E [28] proposes a diffusion model based on point cloud which is flexible but lacks the capability of representing watertight and solid surface of 3D content. To handle this, MeshDiffusion [24] and DiffRF [26] adopt mesh and voxel as the representation. Although explicit representation has superiority in querying and evaluating, those explicit 3D representation methods are memory-expensive when using high-resolution grids. Different from explicit representation, the neural field-based methods [27, 16] adopt an implicit way to encode the geometry and texture information. Despite low memory loading in high-resolution grids, it will cost more time to evaluate. Compared with pure explicit or implicit methods, the hybrid representation triplane achieves a promising trade-off between efficiency and performance [40, 35, 4, 5]. Recently, TRELLIS [48] proposes a powerful 3D generation framework that uses large-scale high-quality 3D data and 2D&3D reference input for reconstruction. Based on it, PhysXGen [2] and PhysX-Anything [3] introduces the first physical-grounded and simulation-ready 3D generative method. Despite introducing multi-modal data, the early fuse framework limits its scalability across different modalities and still suffers from the scarcity of 3D data. To avoid this problem, we integrate the strength of multi-modal data in the later module. In this way, our model not only utilizes the potential of current 3D data to raise the performance but also inherently supports the tokenization of heterogeneous multi-modal inputs through its extensible design.
II-B3 Reconstruction-based methods
@@ III Methods
III-A Collaborative multi-modal coding
As the core component of TriMM, Collaborative Multi-Modal Coding consists of (1) modality-specific encoders and (2) a shared decoder. Specifically, the encoder has three branches, i.e., RGB, RGB-D, and point cloud. These encoders exploit the dense texture information in RGB and the spatial details in depth and point clouds, while the shared decoder ensures that multi-modal inputs are projected into a unified latent space.
Moreover, since our Collaborative Multi-Modal Coding is trained on public 3D datasets, it can be used to alleviate the scarcity of 3D data for generative tasks. By converting heterogeneous RGB images, RGB-D images, and point cloud datasets into a common triplane-based latent space, we can introduce diverse training data and enhance the capacity to model complex 3D structures.
@@ III-A Collaborative multi-modal coding
III-B Triplane latent diffusion model
@@
To validate the modality interaction in our method, we visualize and compare the similarity score of multi-modality triplanes and our generated results in Fig. 5. The results show that the original multi-modality triplanes (top) lack explicit correlation. By integrating these triplanes via diffusion, our generated triplane exhibits clear correlations with all modalities (bottom).
Figure 5: Similarity scores between the multi-modality triplanes and the generated triplane. The results indicate that the original multi-modality triplanes exhibit limited interdependence, whereas our generated triplane effectively integrates and leverages information from all modalities.
@@
IV Experiments
Figure 9: Study of triplanes with different resolutions. To achieve a satisfactory trade-off between efficiency and performance, we adopt the resolution of 96 for our generated results.
Figure 10: Qualitative comparison with text-conditional generation method.
IV-A Implementation details
@@
Training stages In this subsection, we report more details about our TriMM. We compare the parameters, training, and inference time of different modules in Table I. We use 24 NVIDIA A100 GPUs to train our model. Note that the inference time is tested on NVIDIA A100 GPUs. We split our model into two stages: 1) multi-modal coding training including RGB, RGBD, point cloud tokenizer, and shared decoder; 2) Triplane latent diffusion training including VAE and image-conditional diffusion model. For the 1st stage, we first train a RGB reconstruction model similar to other LRM-based methods, e.g., InstantMesh [49]. Besides, to accelerate the convergence of our multi-modal coding, we use the pre-trained RGB triplane to initialize other triplanes. After obtaining all multi-modality triplanes, we train a typical latent diffusion model (2nd stage). Considering the application, we adopt a single image as the condition to train the generative model. We have conducted additional experiments with different triplane resolutions and observed that higher resolutions consistently improve reconstruction performance shown in Fig. 9. However, they also introduce substantially heavier memory consumption and greater difficulty for diffusion training. Therefore, we adopt a resolution of 96×96 in our final model as a balanced trade-off between accuracy and efficiency.
@@
IV-B Representation learning results
@@
IV-C Triplane diffusion results
IV-C1 Qualitative evaluations
In this subsection, we compared our method with our baseline and the most recently existing works including TripoSR [38], CRM [42], LGM [36], InstantMesh [49], and TRELLIS [48]. Baseline represents the generative model trained only on RGB triplanes. We adopt the pre-trained model LGM (trained on 80k data), CRM (trained on 376k data), InstantMesh (trained on 270k data), and TRELLIS (trained on 500k data). We compared our TriMM with other SOTA methods qualitatively in Fig. 7. We can notice that the generated 3D meshes of our TriMM are more stable when facing objects with different appearances and structures. TripoSR can generate a promising texture for 3D assets, especially in the frontal view. However, since the triplane representation has limited capacity to model complex geometric structures, it tends to produce overly flattened 3D shapes. Besides, since LGM, CRM, and InstantMesh are built based on a multi-view diffusion model, the performance of those methods is constrained by the performance of the diffusion model which decreases the final performance of image-to-3D. Relying on high-quality large-scale training data, TRELLIS achieves superior robustness and performance in various 3D assets. Different from integrating 2D and 3D data directly, we introduce the multi-modal coding module based on multi-modal data, thereby raising the robustness when facing complex structures and maintaining the capability of extension. Despite using smaller data, by exploring the potential of current 3D assets, our model achieves a competitive performance compared with TRELLIS. More qualitative results are shown in Fig. 11. Although the main focus of this paper is on the image-to-3D generation task, we additionally extend our experiments to the text-to-3D setting. We use Qwen-Image [44] to generate conditioning images from text prompts and then apply our TriMM framework to produce 3D assets. The comparison with the text-to-3D method 3DTopia [14] demonstrates that our approach achieves impressive generative performance.
@@ IV-D Ablation studies
IV-D1 Analysis about multi-modal data.
@@
IV-D2 Analysis about reconstruction loss.
@@ IV-D3 Analysis of 2D/3D supervision
IV-E Further analysis
IV-E1 Multi-modal data to 3D.
@@ IV-E Further analysis
IV-E2 Introducing multi-modalities for 3D generation
@@
IV-E4 Limitation and future work
Despite the promising results, our method still has several limitations and opens up multiple avenues for future work. First, the triplane representation suffers from inefficient information utilization and is inherently memory-bound when scaling to higher resolutions, which restricts its ability to capture very fine-grained geometric and appearance details. A promising direction is to explore alternative or hybrid 3D representations, such as a combination of triplanes with Gaussian splatting, neural implicit fields, or multi-resolution hierarchical variants. In particular, Gaussian splatting has shown strong advantages in terms of rendering quality and efficiency with compact point-based primitives, but it is less straightforward to encode structured articulation and physical attributes. Designing a hybrid representation that leverages the structured factorization of triplanes and the expressiveness and efficiency of Gaussian splats may offer better trade-offs between expressiveness, memory consumption, and differentiable rendering, thereby further improving reconstruction quality. Second, our current use of multi-modality supervision is constrained by the configuration of existing datasets. Many point cloud datasets provide only geometric information, while most RGB-D datasets are scene-centric rather than object-centric. Although object segmentation masks help localize target objects, they cannot fundamentally compensate for the limited resolution and noise in RGB-D sensing. In future work, we plan to investigate more flexible strategies for exploiting heterogeneous and partially missing modalities, as well as to incorporate richer real-world multi-modality data, which we believe will further enhance the robustness of our model and expand the pool of usable 3D assets.